# Knowledge Adaptation as Posterior Correction

### v1

@@ 3.1 Continual Learning as Posterior Correction
In continual learning, the goal is to quickly update $\boldsymbol{\theta}_{t}$ to $\boldsymbol{\theta}_{t+1}$ when a new loss $\ell_{t+1}$ is available. A popular strategy for this is to regularize the problem. Three kinds of regularization methods are commonly employed: (i) parameter-space regularization, (ii) function or prediction-space regularization, (iii) experience or memory replay. Here, we derive these as special cases of the regularization used for posterior correction used in Eq. 4, denoted by
@@
Here, the KL term performs the parameter-space regularization while the correction term in the data-space is related to the other two regularization methods. Posterior correction naturally combines these seemingly different approaches in one single method.
@@
Next, we consider the function- or prediction- space regularizations. These can be recovered through the correction term. For example, consider a squared-loss
$\ell_{i}(\boldsymbol{\theta})=\mathcal{L}\left[y_{i},\,\hat{y}_{i}(\boldsymbol{\theta})\right]=\mbox{$\frac{1}{2}$}\|y_{i}-\hat{y}_{i}(\boldsymbol{\theta})\|^{2}$
@@
$\begin{split}\ell_{i}(\boldsymbol{\theta})-\hat{\ell}^{\text{iso}}_{i|t}(\boldsymbol{\theta})&=\ell_{i}(\boldsymbol{\theta})-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}[\nabla\ell_{i}]\\ &=\mbox{$\frac{1}{2}$}\|y_{i}-\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}\|^{2}-\boldsymbol{\theta}^{\top}\mbox{$\mbox{$\mathbf{x}$}$}_{i}(\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\mathbf{m}_{t}-y_{i})\\ &=\mbox{$\frac{1}{2}$}\mathcal{L}\left[\hat{y}_{i}(\mathbf{m}_{t}),\,\hat{y}_{i}(\boldsymbol{\theta})\right]+\text{const.}\end{split}$ (7)
This is a function-space regularization used by Benjamin et al. (2019) for neural networks. A visualization of corrections as prediction mismatch is shown in the left panel of Fig. 3. The right panel additionally compares mismatches of two base models. The model with less mismatch is closer to the new model and can facilitate faster adaptation.
Figure 3: Illustration of correction as prediction mismatch for linear regression with $q^{\text{iso}}$ family. The old model is trained on the gray ‘o’ and the new model additionally includes the black ‘$\times$’ too. Corrections are mismatches over old examples (dashed red lines). The right figure shows an additional model (in blue) which has larger mismatches.
In general, this result can be extended to a general nonlinear model $f_{i}(\boldsymbol{\theta})$ (such as neural networks) trained on a loss function $\mathcal{L}(y,\,\hat{y})$ derived using Bregman divergence. As shown in Sec. 4.3, we can write the correction as a sum of a prediction-matching term and an error term arising due to the non-linearity of the model,
$\ell_{i}(\boldsymbol{\theta})-\hat{\ell}^{\text{iso}}_{i|t}(\boldsymbol{\theta})\,\,\approx\,\,\underbrace{\mathcal{L}\left[\hat{y}_{i|t},\,\hat{y}_{i}(\boldsymbol{\theta})\right]}_{\text{Pred. matching}}+r_{i|t}\underbrace{\left[f_{i}(\boldsymbol{\theta})-\hat{f}^{\text{lin}}_{i}(\boldsymbol{\theta})\right]}_{\text{Linearization error}}.$ (8)
Here, we denote by $\hat{y}_{i|t}=\mathbb{E}_{q_{t}}[\hat{y}_{i}(\boldsymbol{\theta})]$ the average posterior predictions under $q_{t}$ and by $r_{i|t}=\hat{y}_{i|t}-y_{i}$ its residual. The term $\hat{f}^{\text{lin}}_{i}(\boldsymbol{\theta})$ is a $1^{\text{st}}$-order linearization of $f_{i}(\boldsymbol{\theta})$ at $\mathbf{m}_{t}$. The second term disappears when the model is linear, leading to a purely mismatch-based correction as in Eq. 7. The approximation is due to the delta method and can be ignored.
The first term favors prediction matching of the past and future, while the second term deals with the residuals obtained because of past mistakes. The expression shows that the mistakes are propagated due to non-linearity of the model. The expression suggests that, for linear cases, prediction matching is sufficient, but otherwise we need additional mechanisms to fix past mistakes. A straightforward approach to reduce this error would be to use a better posterior form. For instance, when using the $q^{\text{full}}$ family, the $\hat{f}_{i}^{\text{lin}}(\boldsymbol{\theta})$ is replaced by a better quadratic surrogate, as shown in Sec. 4.3, thereby reducing the error. In fact, for squared loss, the correction term completely vanishes. The need for correction is reduced with more flexible posterior, which can speed up the adaptation process.
Flexible posteriors however can be costly, for example, computing full covariances is infeasible for larger neural networks. The expression suggests a cheaper alternative which is to simply use replay of past examples to improve continual learning. A similar approach is derived by Daxberger et al. (2023) by using the K-prior framework of Khan and Swaroop (2021), where they combine the three types of regularization methods. We show in Sec. 4.4 that K-prior, which also involves prediction matching, is a special case of posterior correction with $q^{\text{iso}}$ family. Posterior correction generalizes K-prior by allowing the use of an arbitrary posterior distribution. This extends the prediction matching to other kinds of matching, such as, gradient and Hessian matching. It also provides a simpler way to mix and match various regularization methods as discussed in Sec. 4.4.
In Sec. 4.4, we also briefly discuss connections to other approaches covered under the K-prior framework, such as, knowledge distillation (Hinton et al., 2015), incremental Support Vector Machines (Cauwenberghs and Poggio, 2001; Liang and Li, 2009), Similarity Control (Vapnik and Izmailov, 2015), and memory-based continual learning (Lopez-Paz and Ranzato, 2017). Finally, we show an application to sequential Bayesian inference by deriving a method by Bui et al. (2017) for sequential Sparse Variational Gaussian Process as a special instance of posterior correction.
3.2 Influence Estimation as Posterior Correction
Influence estimation aims to estimate the influence of data examples on the models trained over them. For example, given a model $\boldsymbol{\theta}_{t}$ trained on $t$ data examples, we may want to estimates the new model $\boldsymbol{\theta}_{t\backslash j}$ trained on all the data except the $j$’th example. Influence estimates provide a simple expression, often obtained using a Taylor’s approximation evaluated at $\boldsymbol{\theta}_{t}$. These estimates can also be derived as different instances of posterior correction. Here, we will show two results to derive TracIn (Pruthi et al., 2020) and Influence Function (IF) (Jaeckel, 1972; Cook, 1977; Koh and Liang, 2017).
We start by rewriting influence calculation as posterior correction. Taking a variational reformulation, we assume that $q_{t}$ is given and our goal is to estimate the posterior $q_{t\backslash j}$ trained on all example except the $j$’th one. This is shown below in the first line and the following lines are obtained similarly to Eq. 4 to get the correction,
$\begin{split}q_{t\backslash j}&=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[-\ell_{j}]+\sum_{i=1}^{t}\mathbb{E}_{q}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}]\\ &=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[-\ell_{j}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]+\sum_{i=0}^{t}\mathbb{E}_{q}[\ell_{i}-\hat{\ell}_{i|t}]\\ &=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[-\hat{\ell}_{j|t}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]+\sum_{i=0:t,i\neq j}\mathbb{E}_{q}[\ell_{i}-\hat{\ell}_{i|t}]\\ \end{split}$ (9)
@@ 3.2 Influence Estimation as Posterior Correction
The TracIn estimator uses the gradient $\nabla\ell_{j}(\boldsymbol{\theta}_{t})$ as the influence estimator. This follows as a special case of the above by using the $q^{\text{iso}}$ family and solving the following objective where we ignore the correction term,
@@
$\begin{split}&q_{t\backslash j}\approx\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[-\hat{\ell}_{j|t}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]\\ &\implies\mathbf{m}_{t\backslash j}\approx\arg\min_{\mathbf{m}}\,\,\mathbf{m}^{\top}\mathbb{E}_{q_{t}}[-\nabla\ell_{j}]+\mbox{$\frac{1}{2}$}\|\mathbf{m}-\mathbf{m}_{t}\|^{2}\\ &\implies\mathbf{m}_{t\backslash j}\approx\mathbf{m}_{t}+\nabla\ell_{j}(\mathbf{m}_{t})\end{split}$
where the last line is obtained by using the delta method to approximate $\mathbb{E}_{q_{t}}[\nabla\ell_{i}]\approx\nabla\ell_{i}(\mathbf{m}_{t})$. Renaming $\mathbf{m}_{t}$ as $\boldsymbol{\theta}_{t}$, we get the TracIn estimator $\boldsymbol{\theta}_{t\backslash j}-\boldsymbol{\theta}_{t}\approx\nabla\ell_{j}(\boldsymbol{\theta}_{t})$.
@@
Unlike TracIn, the IF uses a Newton step which can be derived by approximating the correction term instead of ignoring it. Therefore it is expected to be more accurate. Specifically, we approximate the correction for all examples from $i=1,2,\ldots,t$ by using the $2^{\text{nd}}$-order Taylor’s expansion of $\ell_{i}$ at $\mathbf{m}_{t}$, as shown below:
@@
$\begin{split}\mathbb{E}_{q}[\ell_{i}-\hat{\ell}_{i|t}]&\approx\mathbb{E}_{q}\left[\boldsymbol{\theta}^{\top}\nabla\ell_{i}(\mathbf{m}_{t})+\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\nabla^{2}\ell_{i}(\mathbf{m}_{t})(\boldsymbol{\theta}-\mathbf{m}_{t})-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}[\nabla\ell_{i}]\right]\\ &\approx\mathbb{E}_{q}\left[\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\nabla^{2}\ell_{i}(\mathbf{m}_{t})(\boldsymbol{\theta}-\mathbf{m}_{t})\right]\\ &=\mbox{$\frac{1}{2}$}(\mathbf{m}-\mathbf{m}_{t})^{\top}\nabla^{2}\ell_{i}(\mathbf{m}_{t})(\mathbf{m}-\mathbf{m}_{t})+\text{const.}\end{split}$ (10)
The second line is obtained again with $\mathbb{E}_{q_{t}}[\nabla\ell_{i}]\approx\nabla\ell_{i}(\mathbf{m}_{t})$ and simplifying. Assuming a quadratic regularizer $\ell_{0}=\mbox{$\frac{1}{2}$}\delta\|\boldsymbol{\theta}\|^{2}$, we can also simplify the sum as shown in Eq. 35. With this, the influence procedure in Eq. 9 simplifies to
@@
$\begin{split}\mathbf{m}_{t\backslash j}&\approx\arg\min_{\mathbf{m}}\,\,-\mathbf{m}^{\top}\mathbb{E}_{q_{t}}[\nabla\ell_{j}]+\mbox{$\frac{1}{2}$}(\mathbf{m}-\mathbf{m}_{t})^{\top}\mathbf{H}_{t\backslash j}(\mathbf{m}-\mathbf{m}_{t})\,\,\approx\mathbf{m}_{t}+\mathbf{H}_{t\backslash j}^{-1}\nabla\ell_{j}(\mathbf{m}_{t})\end{split}$ (11)
@@
where $\mathbf{H}_{t\backslash j}=\sum_{i}\nabla^{2}\ell_{i}(\mathbf{m}_{t})$ where the sum runs over $i\neq j$. This is the IF for removal of an example (Nickl et al., 2023, Eq. 2). The derivation clearly shows the approximations used by TracIn and IF: the former ignores the corrections while the latter reduces it by using Taylor’s approximation. As a result, we expect IF to be a better estimate.
A recent proposal by Nickl et al. (2023) called the Memory-Perturbation Equation generalizes the influence estimation by using the Bayesian Learning Rule. This approach too can be seen as a special instance of posterior correction where a generic exponential family is used but the correction term is ignored. Let us denote the natural parameters of $q_{t}$ and $q_{t\backslash j}$ by $\boldsymbol{\lambda}_{t}$ and $\boldsymbol{\lambda}_{t\backslash j}$, respectively. Then, by using the definition of ^$\hat{\ell}_{t|j}$ from Eq. 3 and ignoring the correction term, we can rewrite Eq. 9 as follows:

### v3

@@ 3.1 Continual Learning as Posterior Correction
In continual learning, the goal is to quickly update $\boldsymbol{\theta}_{t}$ to $\boldsymbol{\theta}_{t+1}$ when a new loss $\ell_{t+1}$ is available. A popular strategy for this is to use regularization. Three kinds of regularization methods are commonly employed: (i) parameter-space regularization, (ii) function or prediction-space regularization, (iii) experience or memory replay. We will show that such regularizers naturally arise as special cases of the regularizer used in Eq. 4 for posterior correction. The regularizer is shown below and denoted by
@@
The KL regularizer operates in the parameter space and the correction term adds further regularization in the data-space, containing as special cases the function and prediction regularization, as well as the experience and memory replay methods. Overall we will show that posterior correction naturally combines all these regularizers into a single regularizer.
3.1.1 Parameter-space regularization as Posterior Correction
We start with the parameter-space regularization which is implemented in $\mathcal{K}_{t}(q)$ through the KL regularizer. Dropping the correction term, simply reduces to Eq. 2 which is known as Variational Continual Learning (cuong2018). Different choices of distribution give rise to different regularizers in the parameter space. For instance, for Gaussian families, the KL term reduces to a quadratic regularizer, shown below for the $q^{\text{iso}}$ and $q^{\text{full}}$ families,
$\mathbb{D}_{\text{KL}}[q^{\text{iso}}\,\|\,q^{\text{iso}}_{t}]$ $=\mbox{$\frac{1}{2}$}(\mathbf{m}-\mathbf{m}_{t})^{\top}(\mathbf{m}-\mathbf{m}_{t})+\text{const.}$ (7)
$\mathbb{D}_{\text{KL}}[q^{\text{full}}\,\|\,q^{\text{full}}_{t}]$ $=\mbox{$\frac{1}{2}$}(\mathbf{m}-\mathbf{m}_{t})^{\top}\mbox{$\mbox{$\mathbf{S}$}$}_{t}(\mathbf{m}-\mathbf{m}_{t})+\mbox{$\frac{1}{2}$}\mbox{Tr}\left[\mbox{$\mbox{$\mathbf{S}$}$}^{-1}\mbox{$\mbox{$\mathbf{S}$}$}_{t}\right]-\mbox{$\frac{1}{2}$}\log\left|\mbox{$\mbox{$\mathbf{S}$}$}^{-1}\mbox{$\mbox{$\mathbf{S}$}$}_{t}\right|+\text{const.}$
The curvature of the regularizer is controlled through the posterior covariance, for instance, using the family $\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m},\mbox{$\mbox{$\mathbf{I}$}$}/\delta)$ with a scalar precision $\delta<0$ will lead to the weighting of the quadratic term by $\delta$. Similarly, for the $q^{\text{full}}$ family, the curvature can be set to the Hessian if we invoke the delta method thereby recovering the so-called Laplace approximation, as detailed in khanRue23. A similar choice with the $q^{\text{diag}}$ family recovers the popular Elastic-Weight Consolidation (kirkpatrick2017overcoming). In general, employing an arbitrary exponential-family, the KL term enables generic parameter-space regularizations as in online learning (hoeven2018many).
3.1.2 Function-space regularization as Posterior Correction
The correction term adds further regularization on top of the KL term, which encompasses function- and prediction- space regularizations. A simple example illustrating this is the squared loss with a linear predictor $\hat{y}_{i}(\boldsymbol{\theta})=\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}$ for a given input-output pair $(\mbox{$\mbox{$\mathbf{x}$}$}_{i},y_{i})$,
$\ell_{i}(\boldsymbol{\theta})=\mathcal{L}\left[y_{i},\,\hat{y}_{i}(\boldsymbol{\theta})\right]=\mbox{$\frac{1}{2}$}\|y_{i}-\hat{y}_{i}(\boldsymbol{\theta})\|^{2}.$
For this loss, the correction term simplifies as ‘prediction matching’ among the new and old model when we use a $q^{\text{iso}}$ family. This is shown below:
$\begin{split}\ell_{i}(\boldsymbol{\theta})-\hat{\ell}^{\text{iso}}_{i|t}(\boldsymbol{\theta})&=\ell_{i}(\boldsymbol{\theta})-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}[\nabla\ell_{i}]\\ &=\mbox{$\frac{1}{2}$}\|y_{i}-\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}\|^{2}-\boldsymbol{\theta}^{\top}\mbox{$\mbox{$\mathbf{x}$}$}_{i}(\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\mathbf{m}_{t}-y_{i})\\ &=\mbox{$\frac{1}{2}$}\|\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\mathbf{m}_{t}-\mbox{$\mbox{$\mathbf{x}$}$}_{i}^{\top}\boldsymbol{\theta}\|^{2}+\text{const.}\end{split}$ (8)
This is equivalent to function-space regularization (benjamin2018measuring). It also corresponds to Knowledge Distillation (hinton2015distilling) under large temperature and centered logits. A visualization is shown in Fig. 4 to highlight the intuition that a model with smaller corrections is also closer to the final goal, and therefore can facilitate faster adaptation.
Figure 4: Visualization of corrections as prediction mismatches for linear regression with $q^{\text{iso}}$ family. The old model is trained on the gray ‘o’ and the new model additionally includes the black $\times$’ too. Corrections are mismatches over old examples (dashed, vertical red lines). The right figure shows an additional slightly-worse model (in blue) with larger mismatches. The red model has smaller correction because it is closer to the black model which also supports the intuition that smaller corrections imply faster adaptation.
3.1.3 Prediction Matching as Posterior Correction of Neural Networks
The result extends to more general cases where corrections take a form similar to Knowledge Distillation, but also account for errors arising due to the model’s non-linearity. To see this, we consider a general nonlinear model such as a neural network, denoted by $f_{i}(\boldsymbol{\theta})$ and trained by penalizing its predictions $\hat{y}_{i}$ through a Bregman loss $\mathcal{L}(y_{i},\,\hat{y}i)$. For example, we can use the cross-entropy loss where $\hat{y}_{i}$ are obtained by passing $f_{i}(\boldsymbol{\theta})$ through the softmax function. We show in Sec. 4.3 that, for such cases, the correction can be written as prediction-matching with an additional error term arising due to non-linearity,
$\ell_{i}(\boldsymbol{\theta})-\hat{\ell}^{\text{iso}}_{i|t}(\boldsymbol{\theta})\,\,=\,\,\underbrace{\mathcal{L}\left[\hat{y}_{i|t},\,\hat{y}_{i}(\boldsymbol{\theta})\right]}_{\text{Pred. matching}}+r_{i|t}\underbrace{\|f_{i}(\boldsymbol{\theta})-\hat{f}^{\text{lin}}_{i}(\boldsymbol{\theta})\|^{2}}_{\text{Linearization error}}+\epsilon_{i|t}.$ (9)
The first term takes a form similar to distillation where we denote by $\hat{y}_{i|t}=\mathbb{E}_{q_{t}}[\hat{y}_{i}(\boldsymbol{\theta})]$ the average posterior predictions of $q_{t}$. The sampling can be seen as playing a role similar to the temperature. In the second term, we denote the residual by $r_{i|t}=\hat{y}_{i|t}-y_{i}$ and the first-order linearization of $f_{i}(\boldsymbol{\theta})$ at $\mathbf{m}_{t}$ by $\hat{f}^{\text{lin}}_{i}(\boldsymbol{\theta})=f_{i}(\mathbf{m}_{t})+(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\nabla f_{i}(\mathbf{m}_{t})$. The linearization error disappears when the model is linear, leading to a purely mismatch-based correction as in Eq. 8. The $\epsilon_{i|t}$ is an error term that arise due to the use of the delta method to approximate the expectation over $q_{t}$ at its mean $\mathbf{m}_{t}$. The error is expected to be small and can be safely ignored.
Overall, the expression shows that the past mistakes are propagated due to non-linearity, and therefore cause no issues at all for linear cases where prediction matching would be sufficient for accurate continual learning. For non-linear models, past mistakes can hurt through examples where the linearization error in $f_{i}(\boldsymbol{\theta})$ is large. For such examples, the correction term suggests to avoid abrupt changes in the feature.
3.1.4 How to reduce corrections and enable accurate continual learning?
A straightforward way to reduce the correction is by using a more accurate posterior. We show this in Sec. 4.3. By using the $q^{\text{full}}$ family, the $\hat{f}_{i}^{\text{lin}}(\boldsymbol{\theta})$ is replaced by a quadratic surrogate $\hat{f}_{i}^{\text{quad}}(\boldsymbol{\theta})=\hat{f}_{i}^{\text{lin}}(\boldsymbol{\theta})+\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\nabla^{2}f_{i}(\mathbf{m}_{t})(\boldsymbol{\theta}-\mathbf{m}_{t})$. A quadratic surrogate yields a smaller linearization error, thereby reducing the harm caused by past mistakes. In fact, for a squared loss, the correction term completely vanishes, giving rise to the classical Bayes’ update for linear regression, where exact continual learning is possible just with the KL term. The example demonstrates that the need for correction is reduced with more flexible posteriors, which can speed up the adaptation process.
Flexible posteriors however can be costly. For example, computing full covariances is infeasible for larger neural networks. For such cases, Eq. 9 suggests a cheaper alternative, which is to simply use replay of past examples to improve continual learning. The split of errors shown in Eq. 9 motivates the use of mixing knowledge distillation with experience replay, as done in Dark Experience Replay (buzzega2020dark). Another approach is derived in daxberger2023improving by using the K-prior framework of khan2021knowledge, where they combine the three types of regularization methods. We show in Sec. 4.4 that the K-prior, which also involves prediction matching, is a special case of posterior correction with the $q^{\text{iso}}$ family. Posterior correction generalizes the K-prior by allowing the use of an arbitrary posterior distribution. This extends the prediction matching to other kinds of matching, such as, gradient and Hessian matching. It also provides a simpler way to mix and match various regularization methods as discussed in Sec. 4.4.
In Sec. 4.4, we also briefly discuss connections to other approaches covered under the K-prior framework, such as incremental Support Vector Machines (cauwenberghs2001incremental; liang2009incremental), Similarity Control (vapnik2015learning), and memory-based continual learning (lopez2017gradient). Finally, we show an application to sequential Bayesian inference by deriving a method by bui2017streaming for sequential Sparse Variational Gaussian Process as a special instance of posterior correction.
3.2 Unlearning as Posterior Correction
@@
[…]
@@
3.2.1 TracIn as Posterior Correction
The TracIn estimator uses the gradient $\nabla\ell_{j}(\boldsymbol{\theta}_{t})$ as the influence estimator. This can be derived as a simplification of Eq. 10 where we make two changes. First, we restrict the set of distributions to the $q^{\text{iso}}$ family, denoted by $\mathcal{Q}^{\text{iso}}$, with candidates $q=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m},\mbox{$\mbox{$\mathbf{I}$}$})$. Second, we ignore the correction term. With these approximations, TracIn is recovered from posterior correction as shown below, where in the first line we make the two changes, then in the second line we rewrite the objective as an optimization problem over $\mathbf{m}$, and in the last line we show the solution:
@@ 3.2.1 TracIn as Posterior Correction
$\begin{split}&q_{t\backslash j}\approx\operatorname*{arg\,min}_{q\in\mathcal{Q}^{\text{iso}}}\,\,\mathbb{E}_{q}[-\hat{\ell}_{j|t}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]\\ &\implies\mathbf{m}_{t\backslash j}\approx\operatorname*{arg\,min}_{\mathbf{m}}\,\,\mathbf{m}^{\top}\mathbb{E}_{q_{t}}[-\nabla\ell_{j}]+\mbox{$\frac{1}{2}$}\|\mathbf{m}-\mathbf{m}_{t}\|^{2}\\ &\implies\mathbf{m}_{t\backslash j}\approx\mathbf{m}_{t}+\nabla\ell_{j}(\mathbf{m}_{t}).\end{split}$ (11)
[…]
3.2.2 Influence Function as Posterior Correction
@@
Unlike TracIn, the IF method uses a Newton step, but this can also be derived through a similar procedure by tightening the approximations made. Specifically, instead of completely ignoring the correction term, we approximate it by using the second-order Taylor expansion
$\ell_{i}(\boldsymbol{\theta})\approx\ell_{i}(\mathbf{m}_{t})+(\boldsymbol{\theta}-\mathbf{m})^{\top}\nabla\ell_{i}(\mathbf{m}_{t})+\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\nabla^{2}\ell_{i}(\mathbf{m}_{t})(\boldsymbol{\theta}-\mathbf{m}_{t}).$
Using this, the correction term is approximated by a quadratic. To simplify the derivation, we assume $\ell_{0}=\mbox{$\frac{1}{2}$}\delta\|\boldsymbol{\theta}\|^{2}$ for a scalar $\delta>0$, although the derivation holds more generally. For this case, we show in App. D, that the correction for the $q^{\text{iso}}$ family simplifies to
[…]
Using this, we can write the correction term as a quadratic,
@@
$\begin{split}\sum_{i=0,i\neq j}^{t}\mathbb{E}_{q}[\ell_{i}-\hat{\ell}_{i|t}]&=\sum_{i=0,i\neq j}^{t}\mathbb{E}_{q}\left[\ell_{i}(\boldsymbol{\theta})-\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}[\nabla\ell_{i}]\right]-\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]+\text{const.}\\ &\approx\sum_{i=0,i\neq j}^{t}\mathbb{E}_{q}\left[\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\nabla^{2}\ell_{i}(\mathbf{m}_{t})(\boldsymbol{\theta}-\mathbf{m}_{t})\right]-\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]+\text{const.}\\ &=\mbox{$\frac{1}{2}$}(\mathbf{m}-\mathbf{m}_{t})^{\top}\mathbf{H}_{t\backslash j}(\mathbf{m}-\mathbf{m}_{t})-\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]+\text{const.}\end{split}$ (13)
The second line is obtained by using the delta method $\mathbb{E}_{q_{t}}[\nabla\ell_{i}]\approx\nabla\ell_{i}(\mathbf{m}_{t})$ which leads to the cancellation of the linear terms, leaving only the quadratic terms. The last line then follows by denoting the Hessian by $\mathbf{H}_{t\backslash j}=\sum_{i=0,i\neq j}^{t}\nabla^{2}\ell_{i}(\mathbf{m}_{t})$.
Using the above approximation, Eq. 10 simplifies to the influence function (IF),
@@
$\begin{split}\mathbf{m}_{t\backslash j}&\approx\operatorname*{arg\,min}_{\mathbf{m}}\,\,\mathbf{m}^{\top}\mathbb{E}_{q_{t}}[-\nabla\ell_{j}]+\mbox{$\frac{1}{2}$}(\mathbf{m}-\mathbf{m}_{t})^{\top}\mathbf{H}_{t\backslash j}(\mathbf{m}-\mathbf{m}_{t})\\ &=\mathbf{m}_{t}+\mathbf{H}_{t\backslash j}^{-1}\mathbb{E}_{q_{t}}[\nabla\ell_{j}]\\ &\approx\mathbf{m}_{t}+\mathbf{H}_{t\backslash j}^{-1}\nabla\ell_{j}(\mathbf{m}_{t}),\end{split}$ (14)
@@
where, similarly to Eq. 11, we use the delta method to approximate the expectation over $q_{t}$. This is the IF for the removal of an example, for example, see nickl2023memory. The derivation clearly shows the approximations used by TracIn and IF: the former ignores the corrections while the latter reduces it by using Taylor’s approximation. IF is a better estimator, at least in theory, because it uses the Hessian as well. This supports our hypothesis that smaller corrections lead to better adaptation methods.
An alternate way to derive the same result is to expand the size of the family from $q^{\text{iso}}$ to $q^{\text{full}}$, which also leads to a reductions in corrections. As shown in App. E, posterior correction in Eq. 10 yields the following update over the $q^{\text{full}}$ family,
[…]
This reduces to Eq. 14 if we use the delta method to approximate $\mathbb{E}_{q_{t}}\left[\nabla\ell_{j}\right]\approx\nabla\ell_{j}(\mathbf{m}_{t})$ and $\mathbb{E}_{q_{t}}\left[\nabla^{2}\ell_{j}\right]\approx\nabla^{2}\ell_{j}(\mathbf{m}_{t})$, and also use $\mbox{$\mbox{$\mathbf{S}$}$}_{t}=\mathbf{H}_{t}$. This shows that we do not have to explicitly approximate the correction (as we did before by using Taylor’s method). We could obtain the same result by simply expanding the size of the posterior class. Both methods lead to similar reductions of correction, and ultimately both lead to the same adaptation methods which is a better estimator than TracIn.
3.2.3 Memory Perturbation as Posterior Correction
[…]